OpenStudy (anonymous):

25x^2-90x=-81 Solve Solving Equations by Factoring

OpenStudy (anonymous):

25x^2-90x+81=0 (5x-9)(5x-9) = 0 (5x-9)=0 5x=9 x=9/5